Ultrasound of intensity 1.70 ✕ 10−2 W/m2 is produced by the rectangular head of a medical imaging device measuring 3.00 cm by 5.25 cm. What is its power output (in W)?
Answer #1

Here,

intensity, I = 1.7 *10^-2 W/m^2

area, A = 3 cm * 5.25 cm

A = 1.575 *10^-3 m^2

for the power output

power output = I * Area

power output = 1.575 *10^-3 * 1.7 *10^-2 W

power output = 2.68 *10^-5 W

the power output is 2.68 *10-5 W